# On the emergence of numerical instabilities in Next Generation Reservoir Computing

### v1

@@ 7. Discussion and conclusions
One of the primary contributors to ill-conditioning is the use of high-degree polynomials. A natural strategy to mitigate this is to select a subset of monomials most relevant to the dynamics. However, identifying the optimal combination of terms is computationally challenging [DMA97]. Greedy algorithms, though suboptimal, have been successfully applied in reservoir computing [DSV+09]. More sophisticated approaches might impose prior structural information, such as sparsity or symmetry, aligning with recent developments in physics-informed machine learning and reservoir computing [CLS21, GGNFN23, GGRG24, BGR+21]. Developing practical algorithms to construct NGRC models that preserve or approximate selected features of the original system remains an open direction.
For complex dynamics, long polynomial expansions (large $p$) may be unavoidable, resulting in severely ill-conditioned matrices and expensive computations. In such cases, one must balance accuracy, interpretability, and computational cost. Recent alternatives address these challenges: for interpretable models, local low-degree polynomial approximations have been proposed [GPB25]; for efficiency, recursive Volterra series expansions offer a delay-agnostic, lower-cost alternative to NGRC [GTO25].
@@
Finally, while this study focused on NGRC, the insights presented here are broadly relevant to other data-driven modeling approaches relying on solving linear equations and least-squares problems. In particular, our findings resonate with similar concerns in Dynamic Mode Decomposition (DMD) [DMM19, PD20], where matrix conditioning also plays a central role. Our results highlight the need for careful hyperparameter selection to ensure numerical stability of NGRC models. This numerical analysis is a first step toward designing better strategies to construct robust reservoir computers.
8. Appendix
8.1. Metrics
@@
Valid prediction time. We use the valid prediction time [VPH+20] to quantify the number of iterations such that the trajectories of the NGRC model and the original dynamics remain close to each other under a certain threshold. For a fixed threshold $\eta=0.9$, we define the valid prediction time as
(22) $\begin{split}n_{VPT}&:=\min_{n}\Big{\{}n:\frac{\|\mathbf{x}_{n}-\mathbf{r}_{n}\|}{\sqrt{\frac{1}{N}\sum_{k=0}^{N}\|\mathbf{x}_{k}\|^{2}}}>\eta\Big{\}}\\ VPT&:=\frac{1}{T_{\Lambda}}n_{VPT}h,\end{split}$
@@
we use B-spline interpolation of degree $3$. Let $\mathcal{I}\subset\mathbb{R}$ be the overlapping domain of both maps, then we define the mean absolute difference over $1000$ evenly spaced points $\{u_{j}\}_{j=1}^{N}\subset\mathcal{I}$ as:
@@
Power spectrum density. To capture the long time statistics of the trajectories we computed the power spectrum density. In the case the trajectories sample the attractor similarly over time, the power spectrum density should be close to other. The attractor similarity is computed using Kullback-Leibler (KL) divergence between the two empirical power spectral densities of the trajectories. More precisely, from the trajectories $\{\mathbf{x}_{n}\}_{n}$ and $\{\mathbf{r}_{n}\}_{n}$ we estimate the power spectrum density $P_{i}$ and $\hat{P}_{i}$ of each $i$th component, respectively, using Welch’s method [Wel67] with Hann window, and compute the divergence error $E$ based on the KL divergence
@@
where $L=\lfloor\frac{5}{dt}\rfloor$ the number of segments, the number of overlapping points $L/2$, computing the periodograms with mean.
@@
Acknowledgments. The authors thank Zheng Bian for enlightening discussions. ERS and EB thank the Collaborative Research in Computational Neuroscience (CRCNS) through R01-AA029926. The ONR, ARO, NIH, and DARPA also support EB.
@@
The code to produce the data that support the findings of this study is available in the https://github.com/edmilson-roque-santos/numerical_stability_NGRC.

### v2

@@ VI Discussion and conclusions
One of the primary contributors to ill-conditioning is the use of high-degree polynomials. A natural strategy to mitigate this is to select a subset of monomials most relevant to the dynamics. However, identifying the optimal combination of terms is computationally challenging [64]. Greedy algorithms, though suboptimal, have been successfully applied in reservoir computing [65]. More sophisticated approaches might impose prior structural information, such as sparsity or symmetry, aligning with recent developments in physics-informed machine learning and reservoir computing [66, 67, 68, 69]. Developing practical algorithms to construct NGRC models that preserve or approximate selected features of the original system remains an open direction. For complex dynamics, long polynomial expansions (large $p$) may remain unavoidable. In such cases, one must balance accuracy, interpretability, and computational cost. Recent alternatives address these challenges. For interpretable models, local low-degree polynomial approximations based on NGRC have been proposed [70]. For efficiency, departing from the equivalence of NGRC and polynomial kernel regression, recursive Volterra series expansions offer a delay-agnostic, lower-cost alternative to NGRC [33].
@@
While this study focused on NGRC, the insights presented here are broadly relevant to other data-driven modeling approaches relying on solving linear equations and least-squares problems. In particular, our findings resonate with similar concerns in Dynamic Mode Decomposition (DMD) [71, 72], where matrix conditioning also plays a central role. Furthermore, the conditioning of the feature matrix itself deserves deeper investigation. Classical numerical analysis states that conditioning is highly sensitive to the distribution of points being evaluated (collocation points or nodes) and the choice of polynomial functions [48]. For example, Vandermonde matrices evaluated at points nearly equally spaced on the unit circle or using orthogonal polynomial bases (like Chebyshev) are known to be well-conditioned. These properties, however, are unrelated to the underlying dynamics. This observation raises an exciting research direction of constructing a basis of orthogonal polynomials adapted to the distribution of points generated by the trajectory of the dynamical system. A first step toward this direction has been applied to coupled maps for network reconstruction [73]. Finally, our results highlight the need for careful hyperparameter selection to ensure numerical stability of NGRC models. This numerical analysis is a first step toward designing better strategies to construct robust reservoir computers.
VII Appendix
VII.1 Metrics
@@
Valid prediction time. We use the valid prediction time [39] to quantify the number of iterations such that the trajectories of the NGRC model and the original dynamics remain close to each other under a certain threshold. For a fixed threshold $\eta=0.9$, we define the valid prediction time as
$\begin{split}n_{VPT}&:=\min_{n}\Big\{n:\frac{\|\mathbf{x}_{n}-\mathbf{r}_{n}\|}{\sqrt{\frac{1}{N}\sum_{k=0}^{N}\|\mathbf{x}_{k}\|^{2}}}>\eta\Big\}\\ VPT&:=\frac{1}{T_{\Lambda}}n_{VPT}h,\end{split}$ (23)
@@
we use B-spline interpolation of degree $3$. Let $\mathcal{I}\subset\mathbb{R}$ be the overlapping domain of both maps, then we define the mean absolute difference over $1000$ evenly spaced points $\{u_{j}\}_{j=1}^{1000}\subset\mathcal{I}$ as:
@@
Power spectrum density. To capture the long time statistics of the trajectories we computed the power spectrum density. In the case the trajectories sample the attractor similarly over time, the power spectrum density should be close to other. The attractor similarity is computed using Kullback-Leibler (KL) divergence between the two empirical power spectrum densities of the trajectories. More precisely, from the trajectories $\{\mathbf{x}_{n}\}_{n}$ and $\{\mathbf{r}_{n}\}_{n}$ we estimate the power spectrum density $P_{i}$ and $\hat{P}_{i}$ of each $i$th coordinate, respectively, using Welch’s method [74] with Hann window, and compute the divergence error $E$ based on the KL divergence
@@
where $L=\lfloor\frac{5}{dt}\rfloor$ is the number of segments, the number of overlapping points $L/2$, computing the periodograms with mean.
Supplementary Material
The Supplementary Material includes a description of additional numerical experiments using the Double-Scroll electronic circuit.
@@
Acknowledgments. The authors thank Zheng Bian and Thomas Peron for enlightening discussions. ERS and EB thank the Collaborative Research in Computational Neuroscience (CRCNS) through R01-AA029926. The ONR, ARO, NIH, and DARPA also support EB.
@@
The code to produce the data supporting this study’s findings is available in the https://github.com/edmilson-roque-santos/numerical_stability_NGRC.
Supplementary Material of
On the emergence of numerical instabilities in Next Generation Reservoir Computing
Edmilson Roque dos Santos, and Erik Bollt.
S1 Double-Scroll electronic circuit
This supplementary material replicates the numerical experiments presented in the main text for the Lorenz-63 system, now applied to the Double-Scroll electronic circuit introduced in [58]. The equations of motion in dimensionless form are given by
$\begin{split}\dot{V}_{1}&=\frac{V_{1}}{R_{1}}-\frac{\Delta V}{R_{2}}-2I_{r}\sinh{(\alpha\Delta V)},\\ \dot{V}_{1}&=\frac{\Delta V}{R_{2}}+2I_{r}\sinh{(\alpha\Delta V)}-I,\\ \dot{I}&=V_{2}-R_{4}I,\end{split}$ (24)
where $\Delta V:=V_{1}-V_{2}$. The parameters are the same used in [10]: $R_{1}=1.2$, $R_{2}=3.44$, $R_{4}=0.193$, $\alpha=11.6$, and $I_{r}=2.25\times 10^{-5}$, resulting in a Lyapunov time of $7.81$-time units. Because the equations in Eq.˜24 include non-polynomial terms, the NGRC model must approximate the original dynamics using polynomial basis functions. Numerical integration uses a fourth-order Runge-Kutta method with a fine time step $0.001$, and the time series is then subsampled at every time step $h$.
Consistent with the main text, we observe: NGRC accurately reproduces the Double-Scroll attractor without requiring regularization. Fig.˜S1 shows the attractor reconstruction and forecasting capability of the NGRC model for the Double-Scroll system. For this example, the condition number $\kappa(\Psi)=1198$ and the closeness of fit for the three coordinates is reported in Table˜3. All algorithms achieve nearly identical closeness of fit up to machine precision. More importantly, the values are much larger than the order $10^{-14}$ as observed for the Lorenz system, primarily due to the $\sinh{}$ term, requiring that polynomial approximation (in this case of degree $3$), inevitably decreasing the fitting accuracy. Since the condition number remains well below the inverse of machine precision, the feature matrix is not in the ill-conditioned regime. As expected, the difference $\Delta$ between algorithms’ solutions is small, on the order of $10^{-12}$, consistent with the discussion in the main text.
Supplementary Figure S1: NGRC reproduces the Double-Scroll electronic circuit attractor. The top panel shows the true (blue) and NGRC reconstructed (red) Double-Scroll attractor. The bottom panel displays NGRC performance during training and testing. The NGRC can forecast the original dynamics up to 5 Lyapunov times. The numerical integrator is the fourth-order Runge-Kutta with $h=0.01$ and the solution is computed by the SVD algorithm. The parameters are delay dimension $k=1$, time lag $\tau=1$, the maximum degree $p=3$, $N_{\mathrm{train}}=10000$, $N_{\mathrm{test}}=80000$ and regularizer parameter $\beta=0$.
Algorithm $\theta_{V_{1}}$ $\theta_{V_{2}}$ $\theta_{I}$
Cholesky $7.765499763260145\times 10^{-2}$ $9.133421534662849\times 10^{-2}$ $2.983845874175234\times 10^{-4}$
SVD $7.765499763260143\times 10^{-2}$ $9.13342153466285\times 10^{-2}$ $2.98384587417524\times 10^{-4}$
LU $7.765499763260143\times 10^{-2}$ $9.13342153466285\times 10^{-2}$ $2.9838458741752435\times 10^{-4}$
Table 3: Closeness of fit for each coordinate of the Double-Scroll system. The parameters are the same as those used for Fig.˜S1.
S1.1 Sparsely-sampled training data
Consider a sparsely-sampled time series scenario for $h=0.25$, similar to the main text, to assess the NGRC performance as the maximum degree $p$ increases, see Fig.˜S2. The findings closely parallel the Lorenz system results: the condition number increases exponentially for increasing $p$, while the closeness of fit $\theta_{\max}$ behaves smoothly only for the SVD algorithm. For Cholesky and LU algorithms, the closeness of fit breaks down at $p=6$, and becomes non-monotonic.
The NGRC performance is quantified by the valid prediction time and the divergence error $E$. NGRC reproduces accurately the attractor without the need for regularization for maximum degrees within the orange band. Unlike the Lorenz system, all three algorithms perform comparably in this regime, with results statistically similar, but degrade for $p>4$. This difference is mostly due to the Double-Scroll system’s non-polynomial vector field. While the Lorenz dynamics can be well-approximated by larger polynomial expansions due to the polynomial structure, the Double-Scroll dynamics is inherently harder to approximate. It is therefore surprising that a third-degree polynomial expansion suffices for useful forecasting.
Supplementary Figure S2: NGRC reproduces the original Double-Scroll dynamics using third-degree polynomial expansions. The top panel shows the maximum closeness of fit over the three components (black) and the condition number (red) with respect to the maximum degree $p$. Subsequent panels show box plots over $25$ initial conditions. The orange shaded area shows the maximum degree where all three algorithms perform equally within statistical significance. The parameters are $h=0.25$, delay dimension $k=1$, time lag $\tau=1$, $N_{\mathrm{train}}=400$, $N_{\mathrm{test}}=6000$ and regularizer parameter $\beta=0$.
S1.2 Partial measurement
Supplementary Figure S3: Methods used for embedding time series inform NGRC modelling of partial measurement of Double Scroll electronic circuit. The top panel shows the closeness of fit $\theta_{V_{1}}$ (black) and condition number of $\Psi$ (red) with respect to the increasing time lag $\tau$. The subsequent panels show the NGRC performance in both metrics, VPT and divergence error $E$. The orange band represents the $\tau=13$, which is informed by the first minimum of mutual information. The blue colormap represents the fraction of bounded NGRC models over the $25$ initial conditions during testing. The dots correspond to the median, and error bars are $25\%$ (lower bound) and $75\%$ (upper bound) quantiles, which are shown only for NGRC models with a fraction larger than $0.5$. The numerical integrator is the fourth-order Runge-Kutta method. The parameters are $h=0.25$, delay dimension $k=3$, maximum degree is $p=3$, $N_{\mathrm{train}}=400$, $N_{\mathrm{test}}=6000$ and regularizer parameter $\beta=0$.
We consider the partial measurement consisting solely of the $V_{1}$-coordinate time series. For $h=0.25$, we choose delay dimension $k=3$ (identified by false nearest neighbors method), based on Fig.˜S2, we set maximum degree $p=3$. Fig.˜S3 shows that by increasing the time lag reduces the condition number monotonically, consistent with the results reported in the main text. The closeness of fit $\theta_{V_{1}}$ exhibits similar behavior as the Lorenz system case, with the difference that, for $\tau>20$, the accuracy decays slightly instead of reaching a plateau.
The NGRC performance is quantified by the valid prediction time, divergence $E$ between the power spectrum densities, and the fraction of bounded NGRC models. Interestingly, the results reveal a wide range of time lags yielding bounded models with statistically similar performance. Although the underlying reason here is unclear, this wide range contrasts with the Lorenz system (Runge–Kutta case), for which there was a narrow range of time lags. Importantly, this range includes the time lag suggested by the first minimum of mutual information method, $\tau=13$, reinforcing that methods used for embedding time series can effectively guide hyperparameter choices. Using these informed hyperparameters, Fig.˜S4 compares the original and reconstructed time series, along with their power spectra. The NGRC model successfully forecasts approximately two Lyapunov times of the original trajectory.
Supplementary Figure S4: NGRC forecasts the partial measurement of the Double-Scroll. The left panel shows the NGRC trajectory (red) compared to the original (blue). The right panel shows the comparison between the power spectrum densities. The parameters are $h=0.25$, delay dimension $k=3$, time lag $\tau=13$, the maximum degree $p=3$, $N_{\mathrm{train}}=400$, $N_{\mathrm{test}}=6000$ and regularizer parameter $\beta=0$.